OpenStudy (anonymous):

Margaret currently has a balance of $7,726.99 in an account earning simple interest. Sixteen years ago she opened the account with an initial deposit of $5,278. What is the interest rate on this account?

OpenStudy (anonymous):

There is the annual compounding equation that looks like FV = PV(1+i)^n where FV = future value [given above] PV = present value [given above] n = number of years [given above] i = interest rate [what we are solving for] make sense now?

OpenStudy (anonymous):

Kinda so it would be 7,726.99=5,278(1+i)^16??

OpenStudy (anonymous):

@Hero

hero (hero):

Looks right. Solve for i

OpenStudy (anonymous):

this is were I get confused.

hero (hero):

Divide both sides by 5,278 then radical both sides to the 16th root.

OpenStudy (anonymous):

so 5,278/7,726.99=0.683060 raise that to the 16th and you get 0.00224566

hero (hero):

I'm pretty sure you did something wrong there.

OpenStudy (anonymous):

I have no idea what im doing wrong I didn't get much practice in on this

OpenStudy (anonymous):

would I multiply the interest by the PV?

hero (hero):

After taking the 16th root of both sides, I get 1.02411 = 1 + i then subtracting 1 from both sides: 1.02411 - 1 = i 1.02411- 1 = i 0.02411 = i

OpenStudy (anonymous):

so the interest would be 2.4%? or would I plug in that to the equation?

hero (hero):

The annual interest rate is 2.4% But for the equation, i = 0.02411

hero (hero):

I remember I was taught to solve these using logs. So dumb when you can just take the nth root of both sides.

OpenStudy (anonymous):

so the new equation would be 7,726.99=5,278(1+0.02411)^16??

hero (hero):

Yes

OpenStudy (anonymous):

I had no time to learn how to do this any other way because I have to take part A and part B this semester it sucks.

hero (hero):

If they make you use logs, you're screwed.

OpenStudy (anonymous):

I remember using logs in other math classes it sucked hopefully I don't have to.

OpenStudy (anonymous):

so part 1 would be 7,726.99=5,278(1.02411)^16th

hero (hero):

I only have one rule: Avoid logs whenever possible.

OpenStudy (anonymous):

then that would become 7,726.99=5,278(1.464)

OpenStudy (anonymous):

after this would I multiply 5278 by 1.464?

hero (hero):

If you're trying to check the result, you can do that, but it's not necessary in this case if you did your algebra correctly. The main thing is making sure the equation used is appropriate for this problem. There are a lot of PV and FV equations out there. It is easy to use one thinking it is the right one just because all the numbers fit and you can solve for the requested variable. But just because you can solve for the requested variable doesn't mean the right equation was used.

OpenStudy (anonymous):

i don't think we used the right one because i did multiply that and got the FV but that percentage is not one of my answers. my answers are 2.0% 9.2% 2.9% 6.3%

hero (hero):

Let me see what happens if I use logs

hero (hero):

If I get a different result with logs, then you have no choice but to learn logs.

OpenStudy (anonymous):

would it be possible to plug in each percent into the equation and try to get the answer that way?

hero (hero):

You could try that, but if the equation is wrong to begin with, it would be useless.

hero (hero):

Also, doing that would work for multiple choice, but if you run into a problem where you have to solve it without multiple choice, you're stuck.

OpenStudy (anonymous):

(1+0.02)^16=1.3727 multiply that by 5278 and you would get 7245 that one is wrong

OpenStudy (anonymous):

2.9 is wrong as well

hero (hero):

There's no way to solve it with logs. Well, actually, there is, if you let b = 1 + i

OpenStudy (anonymous):

i think the formula is \[7,726.99=5,278\( 1+\frac{ R }{ i } )^{t-1}\]

hero (hero):

Where did you get that from? It doesn't look right

OpenStudy (anonymous):

this formula is for compound interest

hero (hero):

The other formula was for compound interest

OpenStudy (anonymous):

A=p(1+r/n)^rt

hero (hero):

I think you mean A = p(1 + r/n)^nt

hero (hero):

It comes out to the exact same thing.

hero (hero):

Because n = 1

hero (hero):

So you would end up with the exact same equation if you used compound interest.

hero (hero):

And by the way, the logs method doesn't work very well.

OpenStudy (anonymous):

damn would it be A=prt+p?

hero (hero):

Wait a minute...there's a future value formula for simple interest.

hero (hero):

The original person who helped you with this used compound interest formula for a simple interest problem.

hero (hero):

Big FAIL

hero (hero):

The correct formula is \(FV = PV(1 + rt)\)

hero (hero):

\(7,726.99=5,278(1+16r)\)

OpenStudy (anonymous):

So it would look like this 7,726.99=5,278(1+r*16)?

OpenStudy (anonymous):

i had it backwards. so then we plug in one of the options right

hero (hero):

And now I know the correct answer.

hero (hero):

I solved for r

hero (hero):

That's what you should do

hero (hero):

It's easy. You don't have to do logs to figure it out

OpenStudy (anonymous):

(1+0.02)16?

hero (hero):

@Shorthell, solve for r

hero (hero):

Trust me, plugging in isn't always the easy way.

hero (hero):

It's a trial and error method at best.

OpenStudy (anonymous):

So how would i solve for r?

hero (hero):

When you solve for r, you will know what the correct answer is.

hero (hero):

First, Divide both sides by 5278 Next, Subtract 1 from both sides Last, Divide both sides by 16

hero (hero):

If you do those steps correctly, you will get the correct answer.

hero (hero):

If you don't, you will not get the correct answer .

OpenStudy (anonymous):

i got 0.022 rounded

hero (hero):

Nope

hero (hero):

You did something wrong

hero (hero):

Don't ask me what you did wrong because you did not post your steps.

OpenStudy (anonymous):

so 7,726.99/5278=1.4638-1=0.4638/16=0.0289 round that and you get 0.029 convert back 2.9%

hero (hero):

:D

OpenStudy (anonymous):

Thank you!

hero (hero):

It's much more satisfying when you solve for r rather than plug and play.

OpenStudy (anonymous):

yes yes it is